Three dice are thrown together. The Probability that the sum of no. appearing on the dice is k whereas the value of k lies between 3 and 8(including 3 and 8)..

Grade:Upto college level

1 Answers

Nishant Vora IIT Patna
9 years ago
Hi Student,
flg are possible cases :
3 = 1+1+1
4= 1+2+1 , 2+1+1, 1+1+2
5 = 1+1+3, 1+3+1, 3+1+1, 1+2+2, 2+1+2, 2+2+1
6 = 1+1+4, 1+4+1, 4+1+1, 1+2+3,2+3+1,3+2+1,2+1+3,3+2+1,…..similarily form all the cases

Then Probablity = 1/6 * 1/6 * 1/6 * number of cases = Ans

Hope this is helpful hint